# ShortGPT: Layers in Large Language Models are More Redundant Than You Expect

### v2

@@ 3.2 Main Results
To validate the efficacy of our proposed method, we conducted comparative experiments against benchmark and baseline techniques commonly employed in large language model evaluation. Considering the current structured pruning methods generally reduce parameters by no more than 30%, we performed experiments with approximately 1/4 of the parameters pruned. The experimental results are presented in Table 1. Additional experiments exploring different parameter reduction proportions will be discussed in the subsequent section.
@@
The results demonstrate that the performance of the model pruned by our method significantly surpasses that of the baseline method, maintaining most of the large language model’s capabilities, such as reasoning, language understanding, knowledge retention, and examination performance, to the greatest extent possible. Furthermore, we note that the approach of reducing the number of layers (ShortGPT/LaCo) outperforms the method of reducing the embedding dimensions (LLMPru./SliceGPT), implying that the model exhibits more redundancy in-depth than in width. Further experimental analysis will be presented in the ensuing section.
4 Analysis
4.1 Varying prune ratio
@@
Figure 5: Pruning by different metrics, with increasing pruning ratio.
@@
As a pruning method, we further validated the effects of different pruning ratios on model performance. Experiments were conducted on the Llama2-7B-Base and Baichuan2-7B-Base models, observing the Perplexity and MMLU. The pruning ratios ranged from 0% to 97%, employing the strategy of deleting layers based on the importance proposed in this paper, resulting in models with 1 to 32 layers. The results in Figure 5 indicate that as the pruning ratio increases, the performance of the model declines. However, the MMLU score drops significantly at a certain layer, which may imply the existence of certain special layers within the network that play a critically important role.
4.2 Depth Redundancy v.s. Width Redundancy
Figure 6: Redundancy on width, removing a single head from model. The x-axis is the head id.
In the preceding sections, we analyzed and explored the redundancy of large language models (LLMs) in terms of depth (layers). However, we also investigated redundancy in terms of width, specifically focusing on the attention heads. Our method involves sequentially removing each head from the multi-head attention mechanism and adjusting the parameters of the attention output projection to ensure the output shape of the Attention block remains unchanged. The Feed-Forward Network (FFN) remains unaltered. Similar to the approach in Section 2.1, we observe the impact of each head’s removal on the final performance of the network. Figure 6 illustrates the changes in Perplexity and MMLU scores after each head is removed. The results indicate that LLMs exhibit a high degree of redundancy in width, comparable to their redundancy in depth. However, this width redundancy does not follow any discernible pattern and varies across different models. We hypothesize that this variation may be attributed to the symmetry among heads.
@@ 4.3 Redundancy on non-transformer LLM
Figure 7: Redundancy on RWKV, removing single layers on RWKV-7B.
We further investigate whether the redundancy observed in depth is a result of the Transformer’s architecture. Given that the vast majority of popular LLMs are based on the Transformer structure. We selected the RWKV-7B model proposed in (Peng et al., 2023), a formidable competitor to the Transformer architecture which indeed consists of identical layers to a certain extent. We analyzed this model using the methodology from 2.1. Figure 7 shows the redundancy of the RWKV-7B 111 We use rwkv-v5-world-7B from https://huggingface.co/BlinkDL/rwkv-5-world model. Through this figure, we can observe that the RWKV-7B model also exhibits a high level of redundancy. This may suggest that the redundancy is universal across LLMs.
4.4 Importance metric
Figure 8: Different importance metrics. Perplexity is calculated by removing each single layer, other metrics is calculated by hidden states of each layer
The overarching concept of our method is to rank layers by importance and remove the less significant ones. The definition of importance metrics has a profound impact on the outcome. In this section, we define and compare various important metrics:
Sequential: The importance is directly proportional to the sequence order, with shallower layers being more important. This can be implemented by taking the negative of each layer’s index as the importance metric.
Reverse-order/Norm: This metric posits that importance is inversely proportional to the sequence order. It assigns higher importance scores to the shallower layers Inspired by Figure 3. This method gives the same order as measuring importance by hidden states norm mentioned in LABEL:layerimportacne. Here, the importance metric is the layer’s index itself.
Relative Magnitude: Proposed in (Samragh et al., 2023), this metric assumes layers with larger $||\frac{f(x)}{x+f(x)}||$ are of higher importance, where $f$ is the layer transformation function.
BI: BI mentioned in previous section 2.2.
Figure 8 demonstrates the different metrics. We observe that shallower layers in the LLM network are more crucial than deeper ones. Figure 5 shows the results of removing layers by different metrics, demonstrating that Our proposed BI outperforms other metrics. The method of (Samragh et al., 2023) is highly competitive, indicating that relative values can also reflect the importance to a certain extent. It is worth noting that the hidden states norm seems to be a good metric when only considering the MMLU benchmark, but the perplexity is relatively poor, which significantly impacts tasks related to text generation. This also indicates the limitations of the current LLMs evaluation methods.
@@ 4.5 Orthogonal to Quantization
In this section, we show that our method is orthogonal to quantization methods. We apply our method to a model  [^2] We take the model from https://huggingface.co/TheBloke/Llama-2-7B-GPTQ quantized by GPTQ algorithm. Table 2 shows that our method is compatible with the quantization-like method, which could reduce the footprint further.
Model Ratio/Layer Perplexity MMLU
Baseline 0%/32 8.4999 37.99
ShortGPT 27.1%/23 42.6951 36.69
@@ 5 Limitation
Although our method demonstrates strong competitiveness compared to current pruning methods, with some benchmarks even suggesting that deleting certain layers has no impact on the final result, this does not imply that layer removal comes without drawbacks. Due to the limitations of benchmarking, current evaluations may not fully capture the impact of layer removal on model performance. Our experiments reveal that the effect of layer removal is significantly more pronounced on generative tasks compared to multiple-choice tasks. On benchmarks such as GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021), removing 25% of the layers often leads to a severe performance drop, with scores approaching zero. Nevertheless, this study proves that models maintain strong semantic understanding and processing capabilities after layer removal. Post-training techniques could potentially further mitigate performance losses, which is an area of ongoing research.
@@ 6 Related works
To reduce the inference cost of large language models and increase their practical applications, there have been many recent works on compressing models, which can be classified into four categories: model pruning, knowledge distillation, quantization, and low-rank factorization.
@@
Model pruning: model pruning (LeCun et al., 1989; Han et al., 2015) is a classic and effective method of reducing model redundancy modules to compress models. The model pruning methods mainly include unstructured pruning and structured pruning. The unstructured pruning simplifies an LLM by removing specific parameters without considering its internal structure, such as SparseGPT (Frantar and Alistarh, 2023) and LoRAPrune (Zhang et al., 2023). However, this method disregards the overall LLM structure, resulting in an irregular sparse model composition. Another more practical approach is structured pruning, GUM(Syed et al., 2023) makes an analysis of several structured pruning methods for decoder-only LLMs. LLM-Pruner (Ma et al., 2024) selectively removes non-critical structures according to gradient information. ShearedLLaMA (Xia et al., 2023) employs targeted structured pruning and dynamic batch loading. LaCo (Yang et al., 2024) used layer merging to compress the model. Compared to the previous method, our method is a simple and efficient structured pruning method.
Knowledge distillation: Knowledge distillation (KD) (Hinton et al., 2015; Gou et al., 2021) is another way of compressing models, where a larger teacher network provides knowledge to a smaller student network. Using the prompt response pairs by LLM APIs to fine-tune small models can get promised results (Li et al., 2022). In addition, when we can access transparent teacher models, the result can be improved more. MiniLLM (Gu et al., 2023) uses reverse Kullback-Leibler divergence to prevent the student model from overestimating the low-probability regions of the teacher distribution. DistlLLM (Ko et al., 2024) uses an adaptive off-policy approach designed to enhance the efficiency in utilizing student-generated outputs. However, compared to model pruning, this method often requires higher computational resources.
@@
Quantization: quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022, 2024) is a widely accepted technique in the field of model compression, which can significantly save the storage and computational costs of deep learning models. Traditional models are generally stored as floating-point numbers, but quantization converts them into integers or other discrete forms. LUT-GEMM (Park et al., 2022) quantifies only weights and optimizes matrix multiplication in LLM using BCQ format. SPQR (Dettmers et al., 2023) identifies and isolates abnormal weights, stores them with higher accuracy, and compresses all other weights into 3-4 bits. Our model pruning method and quantization method are orthogonal, which means quantification based on our pruned model can further compress the model.
Low-rank factorization: low-rank decomposition (Cheng et al., 2017; Povey et al., 2018) is a model compression technique aimed at approximating a given weight matrix by decomposing it into two or more smaller matrices with significantly lower dimensions. TensorGPT (Xu et al., 2023) stores large embeddings in a low-rank tensor format, reducing the space complexity of LLMs and making them available on edge devices. Recently, SliceGPT (Ashkboos et al., 2024) has also implemented structured compression of models based on matrix factorization of hidden states and absorbed the new matrix into the parameters of the origin networks.
@@
Model redundancy: researchers have long noticed the significant redundancy in nonlinear models (Catchpole and Morgan, 1997). In recent years, the transformer model architecture has been widely applied, and researchers have also studied its redundancy. In (Bian et al., 2021), researchers analyzed redundancy in attention mechanisms, in which clear and similar redundancy patterns (cluster structure) are observed among attention heads. In (Dalvi et al., 2020), researchers dissect two pre-trained models, BERT (Devlin et al., 2018) and XLNet (Yang et al., 2019), studying how much redundancy they exhibit at a representation level and a more fine-grained neuron-level. However, the redundancy in current large language models based on decoder-only structures still needs to be explored.
@@ 7 Conclusion
This work has introduced a novel approach to pruning Large Language Models (LLMs) based on layer redundancy and an "importance" metric defined as attention entropy. Our study reveals a significant degree of layer-wise redundancy in LLMs, indicating that certain layers contribute minimally to the overall network function and can thus be removed without substantially compromising model performance. By employing a straightforward layer removal strategy guided by the calculated importance of each layer, we have demonstrated that it is possible to maintain up to 95% of an LLM’s performance while reducing the model’s parameter count and computational requirements by approximately 25%. This achievement not only surpasses previous state-of-the-art pruning methods but also underscores the potential for further optimizations in LLM deployment strategies.
Our findings suggest that the redundancy inherent in LLMs is largely depth-based rather than width-based, highlighting an avenue for future research into the structural efficiency of neural networks. Moreover, our pruning approach, characterized by its simplicity and effectiveness, is compatible with other compression techniques such as quantization, offering a composite pathway to model size reduction that is both significant and versatile.
The implications of our research extend beyond academic interest, offering practical solutions for deploying advanced LLMs in resource-constrained environments. By enabling more efficient model architectures without the need for extensive retraining, our pruning method facilitates broader access to cutting-edge AI capabilities across a variety of platforms and devices.
In conclusion, our investigation into layer redundancy and the development of an importance-based pruning strategy represent a meaningful advancement in the optimization of Large Language Models. As the demand for sophisticated AI tools continues to grow, approaches like ours will play a crucial role in making these technologies more accessible and sustainable. Future work will focus on refining our understanding of model redundancy and exploring additional methods for enhancing the efficiency of neural network models.
Appendix A Detailed Strategy for Layer Removal
Table 3: Strategies for Layer Removal in Models.
Strategy
Description
Sequential
Layers are removed sequentially from the beginning of the model. The process starts with layer 0 and progressively includes more layers for removal (e.g., {0}, {0, 1}, …).
Reverse-order
This strategy involves starting from the model’s final layer and progressively removing layers in reverse order (e.g., {-1}, {-1, -2}, …).
Relative Magnitude
Layers are removed in ascending order based on their Relative Magnitude values. The removal process accumulates layers from those with the smallest to the largest values, mirroring the sequential strategy’s accumulation method.
BI (Block Influence)
Follows a similar accumulation approach as the Sequential strategy, but layers are ordered and removed according to their BI values, starting from the lowest and moving to the highest.
Appendix B Benchmark Setup for Removed Layers
Before delineating the specific layer removal configurations in the benchmarked models, a concise overview of the structural modifications is warranted. The Llama-2-7B and Baichuan-2-7B models, initially designed with 32 layers, have been streamlined by excising 9 layers, corresponding to a 28% reduction of the total layer framework. The 13B variants, originally constituted of 40 layers, experienced a decrement of 10 layers, marking a 25% diminution. The selection criterion for layer removal was based on ascending Block Influence (BI) values, ensuring a data-driven approach to minimization. This premise underpins the benchmark configurations.
Table 4: Setup of Removed Layers for Benchmark Models

### v3

@@ 4.2 Main Results
To validate the efficacy of our proposed method, we conducted comparative experiments against baseline techniques commonly employed in large language model evaluation. Considering the current structured pruning methods generally reduce parameters by no more than 30%, we performed experiments with approximately 1/4 of the parameters pruned. The experimental results are presented in Table 2. Additional experiments exploring different parameter reduction proportions will be discussed in the subsequent section.
@@
The results demonstrate that the performance of the model pruned by our method significantly surpasses that of the baseline methods, maintaining most of the large language model’s capabilities. Furthermore, we note that the approach of reducing the number of layers (ShortGPT/LaCo) outperforms the method of reducing the embedding dimensions (LLMPru./SliceGPT), implying that the model exhibits more redundancy in depth than in width. Further experimental analysis will be presented in the ensuing section.
In Table 2, we fully adopted the benchmark, model, and pruning ratio in the LaCo paper. In order to make a more fair comparison with LLMprun. and SliceGPT, we compared them with the same benchmark, model, and pruning ratio in their original paper. The experimental results are shown in Appendix C. Consistent with our findings in Table 2, these experiments further demonstrate the significant layer redundancy present in existing large language models, and ShortGPT achieves superior performance compared to other pruning methods.
The results show that coarse-grained pruning methods, such as removing entire layers, often outperform fine-grained approaches like Slice GPT or LLM Pruner. We speculate that the reason is that the large language model is actually very robust, as shown in Figure 1, removing any deep layer individually actually has very little impact on the final output, which means it is difficult to define the importance of a finer grained module and perform pruning.
4.3 Varying metric and pruning ratio
Figure 4: Comparison of different importance metrics. Perplexity is calculated by removing each single layer, other metrics is calculated by hidden states of each layer.
The core principle of our method is to rank layers by their importance and remove the less significant ones. The choice of importance metric significantly influences the outcome. In this section, we define and compare several different importance metrics:
  • Sequential: The importance is directly proportional to the sequence order, with shallower layers being less important. This can be implemented by assigning the negative value of each layer’s index as its importance metric.
  • Norm/Reverse-order: This metric posits that importance is inversely proportional to the sequence order. It assigns higher importance scores to the shallower layers. This method gives the same order as measuring importance by hidden states norm as Figure 4 shows.
  • Relative Magnitude: Proposed in Samragh et al. (2023), this metric assumes layers with larger $||\frac{f(x)}{x+f(x)}||$ are of higher importance, where $f$ is the layer transformation function.
  • BI: we calculate the BI score mentioned in Section 3.1 as importance metric.
Figure 4 demonstrates the different metrics. We observe that shallower layers in the LLM network are more crucial than deeper ones. Figure 5 shows the results of removing layers by different metrics, demonstrating that Our proposed BI outperforms other metrics. The method of Relative Magnitude is highly competitive, indicating that relative values can also reflect the importance to some extent. It is worth noting that the hidden states norm seems to be a good metric when only considering the MMLU benchmark, but the perplexity is relatively poor.
@@ 4.3 Varying metric and pruning ratio
Figure 5: Performance of MMLU and perplexity when we prune by different metrics, with increasing pruning ratio. We can see that as the pruning ratio increases, the performance of the model declines.
@@
As a pruning method, we further validated the effects of different pruning ratios on model performance. Experiments were conducted on the Llama2 and Baichuan2 models, observing the Perplexity and MMLU. The results for Llama2, as shown in Figure 5, demonstrate that the model’s performance generally declines as the pruning ratio increases. However, we observe a notable phenomenon: the MMLU score exhibits a sharp drop at a specific layer. This sudden decrease suggests the presence of certain critical layers within the network that play a particularly important role in maintaining performance. Similar patterns are observed in the Baichuan2 model, as illustrated in Appendix B.
@@ 4.4 Redundancy on non-transformer LLM
To determine whether the observed depth redundancy is specific to the Transformer architecture, we extended our investigation to include two popular non-Transformer models, RWKV-7B111 We use rwkv-v5-world-7B from https://huggingface.co/RWKV/v5-Eagle-7B-HF (Peng et al., 2023) and Mamba-2.8B  [^2] We take the model from https://huggingface.co/state-spaces/mamba-2.8b-hf (Gu & Dao, 2023). Our experiments revealed that these models also exhibit resilience to layer removal, maintaining performance despite the elimination of certain layers. This finding suggests that the redundancy phenomenon may not be unique to Transformer-based models, but rather a common characteristic across current large language models. Table 3 shows that our method is applicable and effective for both Mamba and RWKV models, suggesting that the redundancy is universal across current LLMs. However, it is worth noting that the RWKV model appears less redundant than Mamba and Transformer models, which warrants further investigation.
Table 3: ShortGPT pruning on RWKV and Mamba.
Model Pruning ratio CMNLI HeSw PIQA CHID WSC CoQA BoolQ Race-H Race-M XSum C3 MMLU CMMLU Ave. Per.
Mamba-2.8B 0% 35.97 61.84 75.52 35.56 49.69 56.35 60.67 24.9 25.3 15.03 42.08 26.29 25.32 41.12 100.00
10.9% 32.95 59.71 73.01 32.52 49.28 52.66 51.41 24.27 25.21 14.95 41.1 26.01 25.00 39.08 95.04
20.3% 31.29 55.69 69.64 29.12 48.36 48.32 62.2 23.61 23.61 14.71 41.59 25.69 25.37 38.36 93.29
25% 29.96 52.38 68.77 26.02 48.26 44.96 62.2 23.67 23.26 14.00 40.71 24.32 24.89 37.18 90.42
31.3% 28.25 47.02 64.91 21.38 49.69 44.96 62.17 21.87 22.77 13.77 40.44 24.48 24.77 35.59 86.55
RWKV-7B 0% 32.07 65.98 77.09 85.36 50.00 62.65 62.72 38.56 45.47 16.5 57.97 31.85 28.54 50.37 100.00
9.4% 32.6 56.41 73.94 78.12 50.00 49.55 62.35 25.9 25.77 9.57 54.68 27.29 25.03 43.94 87.23
18.8% 32.11 49.47 71.55 65.63 50.00 40.54 61.19 22.04 23.75 8.13 49.15 26.35 25 40.38 80.17
25% 32.41 39.73 65.13 52.6 50.00 29.65 60.92 22.56 21.59 12.02 41.86 25.52 25.08 36.85 73.16
28.1% 33.11 32.22 60.01 32.47 50.1 28.34 60.85 22.27 21.31 10.43 37.81 25.64 25.15 33.82 67.14
@@ 4.5 Orthogonal to Quantization
In this section, we show that our method is orthogonal to quantization methods. We apply our method to Llama2-7B  [^3] We take the model from https://huggingface.co/TheBloke/Llama-2-7B-GPTQ quantized by GPTQ algorithm. Table 4 shows that our method is compatible with the quantization-like method. In addition, we compared the performance of applying pruning before quantization  [^4] We use GPTQ algorithm for quantization from https://github.com/AutoGPTQ/AutoGPTQ. The results shown in the Table 5 further indicates that quantization and ShortGPT are orthogonal operations.
@@
Model Ratio/Layer Perplexity MMLU Throughput (speed up)
Baseline 0%/32 8.03 43.17 4331.23 Token/s (1.00x)
3.1%/31 8.37 42.88 4399.31 Token/s (1.02x)
9.4%/29 9.44 42.31 4602.26 Token/s (1.06x)
ShortGPT 12.5%/28 10.24 41.62 4680.68 Token/s (1.08x)
15.6%/27 11.42 43.17 4756.94 Token/s (1.10x)
25.0%/24 22.29 41.68 5045.59 Token/s (1.16x)
27.1%/23 40.78 43.35 5146.99 Token/s (1.19x)
Table 5: Performance comparison of different methods
Method MMLU CMMLU
Llama2-7B-Baseline 45.4 32.9
4-bit quantization 44.9 32.5
Layer removal (27.1%) 44.0 32.3
4-bit quantization then layer removal 42.4 31.0
Layer removal then 4-bit quantization 41.2 30.5
4.6 Post training to restore performance
To mitigate the performance loss resulting from layer removal, we explored post-training strategies inspired by Chen et al. (2024). Our approach comprised two key steps: 1)Replacement: We substituted the removed layers with lightweight Multi-Layer Perceptron (MLP) modules. 2)Retraining: We subsequently retrained the modified model. The results in Table 6 demonstrate the potential of post-train in recover performance loss. Appendix F list the training details.
Table 6: Post-train Llama2-7B to restore performance.
Method Avg. Ratio CMNLI HeSw PIQA CHID WSC CoQA BoolQ Race-H Race-M XSum C3 MMLU CMMLU
Dense 47.78 0% 32.99 71.26 77.91 41.66 50.00 64.62 71.62 35.71 34.19 19.40 43.56 45.39 32.92
ShortGPT 41.22 27.1% 32.95 53.02 66.43 24.68 52.46 47.99 74.41 32.25 35.17 0.67 39.62 43.96 32.25
ShortGPT+post-train 43.16 24.0% 32.99 54.83 68.12 31.82 51.37 58.32 72.36 34.18 34.68 4.89 40.37 44.47 32.73
@@ 5 Limitation
Although our method demonstrates strong competitiveness compared to current pruning methods, there are some phenomena that have not been explained. Our experiments reveal that the negative effect of layer removal is more significant on generative tasks compared to multiple-choice tasks. When we remove 25% layers from Llama2-7B or Baichuan2-7B, the performance in generative tasks such as XSum and C3 deceases to nearly zero, although the performance decline was not as significant on the larger model of the 13B. We speculate that compared to multiple-choice tasks, generative tasks face the problem of accumulated errors and large model is more robust than small one. The reasons behind it still need to be explored. The post-training techniques discussed in Section 4.6 have the potential to mitigate this issue and warrant further exploration.
@@ 6 Related works
To reduce the inference cost of large language models and increase their practical applications, there have been many recent works on compressing models, which can be classified into two categories: model pruning and quantization. Besides, there are some works aim to study the redundancy of model which is essential for compressing models.
@@
Model pruning: model pruning (LeCun et al., 1989; Han et al., 2015) is a classic and effective method of reducing model redundancy modules to compress models. The model pruning methods mainly include unstructured pruning and structured pruning. The unstructured pruning simplifies an LLM by removing specific parameters without considering its internal structure, such as SparseGPT (Frantar & Alistarh, 2023) and LoRAPrune (Zhang et al., 2023). However, this method disregards the overall LLM structure, resulting in an irregular sparse model composition. Another more practical approach is structured pruning, GUM(Syed et al., 2023) makes an analysis of several structured pruning methods for decoder-only LLMs. LLM-Pruner (Ma et al., 2024) selectively removes non-critical structures according to gradient information. ShearedLLaMA (Xia et al., 2023) employs targeted structured pruning and dynamic batch loading. LaCo (Yang et al., 2024) used layer merging to compress the model. Compared to the previous method, our method is a simple and efficient structured pruning method.
@@
Quantization: quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022; 2024) is a widely accepted technique in the field of model compression, which can significantly save the storage and computational costs of deep learning models. Traditional models are generally stored as floating-point numbers, but quantization converts them into integers or other discrete forms. LUT-GEMM (Park et al., 2022) quantifies only weights and optimizes matrix multiplication in LLM using BCQ format. SPQR (Dettmers et al., 2023) identifies and isolates abnormal weights, stores them with higher accuracy, and compresses all other weights into 3-4 bits. Our model pruning method and quantization method are orthogonal, which means quantification based on our pruned model can further compress the model.
@@
Model redundancy: researchers have long noticed the significant redundancy in nonlinear models (Catchpole & Morgan, 1997). In recent years, the transformer model architecture has been widely applied, and researchers have also studied its redundancy. In (Bian et al., 2021), researchers analyzed redundancy in attention mechanisms, in which clear and similar redundancy patterns (cluster structure) are observed among attention heads. In (Dalvi et al., 2020), researchers dissect two pre-trained models, BERT (Devlin et al., 2018) and XLNet (Yang et al., 2019), studying how much redundancy they exhibit at a representation level and a more fine-grained neuron-level. However, the redundancy in current large language models based on decoder-only structures still needs to be explored.
@@ 7 Conclusion
In this work, we uncovered the significant layer-wise redundancy of LLMs, Our research demonstrates that certain layers contribute minimally to overall network functionality and can be removed without substantially compromising model performance. Based on our observation, We introduce Block influence to quantify the importance of each layer and propose a simple and straightforward pruning method: layer removal. Our experiments demonstrates that it is possible to maintain up to approximately 90% of a LLM’s performance while reducing the model’s parameter amount and computational requirements by approximately 25%. Besides, our method is orthogonal to quantization methods and can be further improved by continual training. We hope that our work can provide some insight for future model compression techniques. Moreover, our work suggests potential avenues for improving the efficiency of model training by reducing inherent redundancy in the future.
Appendix A Mathematical explanation for why pre-norm brings high similarity
We provide a simple explanation here about how pre-norm leads to high deep similarity in this section, here we adopt RMSNorm (Zhang & Sennrich, 2019) for convenient, which is also the popular pre-norm used in many recent LLMs, such as Llama and Mamba.
Lemma 1
(Xiong et al., 2020) At initialization, for the Pre-LN Transformer, $(1+\frac{L}{2})d\leq\mathbb{E}(||x_{L,i}||^{2}_{2})\leq(1+\frac{3L}{2})d$ for all $L>0$ and $i$. Expectations are taken over the input and the randomness of initialization, where the hidden state of $L^{th}$ layer is $x_{L}$.
From Lemma 1, the hidden state of the pre-norm model will continuously increase as the number of layers increases. And under the assumption of each component of $x_{l}$ has a mean of 00, we can obtain $||x_{L}||=\Theta(\sqrt{L})$.
Then we consider $x_{L+1}=x_{L}+f_{L}(x_{L},\theta_{L})$, where $f_{L}$ is a operation such as Attention or MLP, $\theta_{L}$ is learnable parameters. Then $f_{L}(x_{L},\theta_{L})=O(1)$ respect to $L$, for Attention as example, $||f_{L}(x_{L},\theta_{L})||=||(softmax(Q^{T}K)X_{L}/||X_{L}||\cdot(\sigma_{rms}))W_{v}W_{q}||=O(||\sigma_{rms}||||W_{v}||||W_{o}||)=O(1)$ respect to $L$.
Then we can get:
cos $\textit{cos similarity}(X_{L+1},X_{L})$ $=\frac{x_{L+1}x_{L}}{||x_{L+1}||||x_{L}||}=\frac{||x_{L}||^{2}}{||x_{L+1}||||x_{L}||}+\frac{f_{L}(x_{L},\theta)x_{L}}{||x_{L+1}||||x_{L}||}$ (2)
$\geq\frac{||x_{L}||^{2}}{||x_{L+1}||||x_{L}||}-\frac{||f_{L}(x_{L},\theta)||||x_{L}||}{||x_{L+1}||||x_{L}||}$ (3)
$=\frac{||x_{L}||}{||x_{L+1}||}-\frac{||f_{L}(x_{L},\theta)||}{x_{L+1}}=\Theta(\sqrt{\frac{L}{L+1})}-O(\sqrt{\frac{1}{L+1}})$ (4)
This means that as the number of layers $L$ increases, the similarity between the input and output of the layer will be high. This means that the role of $f_{L}$ may be relatively small, and removing it from the network may have a relatively small impact to the model.
Although the above theoretical analysis is only for randomly initialized models, this phenomenon that deep layer has similar input and output exists in both our own trained models shown in Figure 2 and existing models in Figure 4.
Appendix B Layer Removal on Baichuan2-series Model
Figure 6: Pruning by different metrics on Baichuan2-series model.
Appendix C A Fair comparison with SliceGPT and LLMprun.
In Table 2, we fully adopted the benchmark, model, and pruning ratio in the LaCo’s paper. For a fair comparison with LLM pruner and SliceGPT, we do the same experiments in the original paper of LLM pruner and SliceGPT. The results is provided in Table 7 and Table 8. We take the same benchmarks, models and pruning ratio as the corresponding original paper. The results demonstrate that our method is highly competitive.
[b]
Table 7: Comparison between ShortGPT and LLM-pruner. The Table is corresponding to the Table 1 of LLM pruner(Zhang et al., 2023).
Model Pruning ratio Method BoolQ PIQA Hellaswag Winogrande Arc-e Arc-c OBQA Avg.
Llama-7B Ratio=0% Baseline 73.18 78.35 72.99 67.01 67.45 41.38 42.4 63.25
Ratio=20% LLM-pruner 59.39 75.57 65.34 61.33 59.18 37.12 39.80 56.82
Ratio=21.9 % ShortGPT 68.26 72.28 61.7 63.77 60.22 39 41.6 58.12
Llama-13B Ratio=0% Baseline 68.47 78.89 76.24 70.09 74.58 44.54 42.00 64.97
Ratio=20% LLM-pruner 67.68 77.15 73.41 65.11 68.35 38.4 42.4 61.79
Ratio=20% ShortGPT 68.41 76.36 72.9 67.4 68.62 39.2 41 61.98
[b] Model Pruning ratio Method PIQA Hellaswag Winogrande Arc-e Arc-c Avg. Llama-2-7B 0% Baseline 79.11 75.99 69.06 74.58 46.25 69 20% SliceGPT 71.87 58.1 63.04 69.87 43.09 63.45 25% SliceGPT 68.55 58.1 62.04 57.46 35.07 56.15 30% SliceGPT 66.1 52.69 56.82 35.07 56.82 56.15 21.9% ShortGPT 72.76 66.39 66.27 59.39 39.85 60.93 25% ShortGPT 70.53 62.68 64.7 58.39 39.51 59.16 31.6% ShortGPT 67.87 62.19 64.38 56.57 40.86 58.37 Llama-2-13B 0% Baseline 80.47 79.39 72.22 77.48 49.23 71.76 20% SliceGPT 71.87 69.38 63.04 69.87 43.09 63.45 25% SliceGPT 68.55 67.48 58.1 62.5 37.88 58.9 30% SliceGPT 66.1 65.11 52.69 56.82 35.07 55.16 20% ShortGPT 76.95 74.67 71.14 69.56 45.63 67.59 25% ShortGPT 74.39 71.65 70.98 67.09 43.93 65.61 30% ShortGPT 72.11 71.93 67.19 61.09 40.88 62.64 Llama-2-70B 0% Baseline 82.7 83.84 77.98 80.98 57.34 76.57 20% SliceGPT 76.61 72.98 74.92 80.51 55.2 72.34 25% SliceGPT 74.92 68.74 74.92 77.9 51.71 69.75 30% SliceGPT 72.31 63.69 73.4 51.71 47.61 66.11 20% ShortGPT 76.02 78.87 71.69 76.02 52.95 71.68 25% ShortGPT 73.2 76.72 71.85 73.2 49.9 69.79 30% ShortGPT 74.44 75.31 72.33 74.44 49.22 69.4
Table 8: Comparison between ShortGPT and SliceGPT. The Table is corresponding to the Table 7 of SliceGPT(Ashkboos et al., 2024).
Appendix D Detailed Strategies for Layer Removal
We list the details of different layer removal strategies in Table 10. The concrete removed layers by ShortGPT in Table 2 are listed in Table 9
Table 9: Setup of Removed Layers for Benchmark Models.